KBL Inc. with $20m funds available is seeking covered interest arbitrage (CIA) in Denmark Spot Rate = 6.15 (Kr/$) 3-month forward rate = 6.20 (Kr/S) 3-Month interest rate in US = 2.75% 3-month interest rate in Denmark = 3.25% The CIA profit potential is: Select one: a. -3.23% b. 0.81% c. -2.73% d. None of the Above e. 0.50%
Answer #1

KBL Inc is having $ 20 Mn, the same would be invested in Danish market by buying Danish Krone in the spot market.

Spot Rate =6.15 (Kr/$)

Danish Krone that can be purchased with 20M $ = 20M*6.15 =123 Mn Kr.

Interest earned in Kr = 123Mn*3.25% / 4 = 999375.

Total Danish Krone at the end of 3 mnths = 123000000+999375 =123999375.

No of $ we get by converting Kr at forward rate of 6.20 (Kr/$) = 123999375/6.20 = $ 19999899.19.

As we can see we ended up with lesser $ resulting in a loss. LOSS = 20000000-19999899.19 = $ 100.81

But the actual loss is higher as KBL Inc could have invested in American market @ 2.75% PA.

Interest in American Market = 20000000*2.75% / 4. =137500$.

Total Loss(Including opportunity Cost) = 137600.81 $ = 0.688% ( Ans is D. None of the above).

There is an Arbitrage potential in the currency rate, We should've borrowed the funds in Denmark and Invested in American market to  make a profit.